OpenStudy (wondermath):

How many solutions does cos2x=sinx+cosx have? Are you supposed to use the double angle formula?

OpenStudy (anonymous):

\[\cos2x=\cos(x+x)=\cos x cosx-sinxsinx=\cos^2x-\sin^2x\] So \[\cos^2x-cosx=sinx+\sin^2x\] No, try again: \[\cos^2x-\sin^2x=(cosx+sinx)(cosx-sinx)\]\[=cosx+sinx\] so cosx - sinx = 1 But this is always true, so infinitely many values of x between zero degrees and 90 degrees. I think!

OpenStudy (agreene):

There are going to be infinite solutions because the left side alternates between 1 and -1... with the right alternating between ~1.5 and -1.5...when they cross is a bit harder to pin down--but there will be infinite solutions.